### v2

# Enabling Physical Localization of Uncooperative Cellular Devices

Abstract
In cellular networks, it can become necessary for authorities to physically locate user devices for tracking criminals or illegal devices. While cellular operators can provide authorities with cell information the device is camping on, fine-grained localization is still required. Therefore, the authorized agents trace the device by monitoring its uplink signals. However, tracking the uplink signal source without its cooperation is challenging even for operators and authorities. Particularly, three challenges remain for fine-grained localization: i) localization works only if devices generate enough uplink traffic reliably over time, ii) the target device might generate its uplink traffic with significantly low power, and iii) cellular repeater may add too much noise to true uplink signals. While these challenges present practical hurdles for localization, they have been overlooked in prior works.
In this work, we investigate the impact of these real-world challenges on cellular localization and propose an Uncooperative Multiangulation Attack (UMA) that addresses these challenges. UMA can 1) force a target device to transmit traffic continuously, 2) boost the target’s signal strength to the maximum, and 3) uniquely distinguish traffic from the target and the repeaters. Notably, the UMA technique works without privilege on cellular operators or user devices, which makes it operate on any LTE network. Our evaluations show that UMA effectively resolves the challenges in real-world environments when devices are not cooperative for localization. Our approach exploits the current cellular design vulnerabilities, which we have responsibly disclosed to GSMA.
1 Introduction
The inherent broadcast nature of wireless signals enables seamless communication without the need for revealing precise physical locations. This very feature, however, poses a significant challenge for authorities with legitimate reasons to pinpoint the exact location of targeted user devices. For example, the perpetrators using SIM boxes engaged in activities like bypass fraud [42, 44] and scams [15, 26] can operate in densely populated urban environments, making it difficult for law enforcement agencies to track and apprehend them. In cases involving criminal activities, it becomes imperative to locate and arrest individuals promptly based on the signals emitted by their cellular devices.
Authorities may seek assistance from mobile operators in the process of cellular device localization, and operators are generally willing to cooperate if the requests are valid and comply with legal procedures [18]. However, mobile operators’ assistance in localization comes with certain limitations. While they can provide authorities with readily-available, high-level user data, such as serving cell ID or triangulation based on signal strength, the level of precision achieved remains relatively coarse-grained. For example, Cell ID (CID) provides only limited positioning of devices with about 1km accuracy [13, 17, 41]. While existing native 3GPP positioning features, such as Observed Time Difference of Arrival (OTDoA) via Positioning Reference Signal (PRS), Uplink Time Difference of Arrival (UTDoA), E-CID (Enhanced CID), have been proposed to improve cellular localization, they are all optional features [3] and thus cannot provide universal fine-grained localization across all mobile operators, radio-access network (RAN) vendors, and User Equipment (UE) manufacturers.
For fine-grained LTE localization localizers require active, real-time monitoring of UE uplink signals, the true signal sources of targeted UEs. This requires additional physical-layer approaches beyond collecting control-plane data in mobile networks; like prior works that follow. For example, Timing Advance (TA) commands used in cellular networks account for the propagation delay of radio signals between the base station and a UE and it can be used to infer sub-cell-level UE locations [48]. Localizers may also determine the distance to a target user by utilizing Time of Arrival (ToA) measurements; e.g., LTrack [31] exploits the ToA of UE uplink signals and the TA commands in the LTE systems. LTEye [32] measures the Angle of Arrival (AoA) of UE uplink signals using a Synthetic Aperture Radar (SAR).
Unfortunately, however, whether these physical-layer cellular localization techniques can provide reliable, fine-grained localization of cellular devices in practice has not been thoroughly tested and demonstrated. Thus, the effectiveness of these techniques under realistic scenarios within a universal cellular infrastructure remains unclear. As a result, this work aims to a) reveal any overlooked challenges of cellular localization in real-world environments, b) address them in a universal setting, and c) demonstrate an end-to-end fine-grained cellular localization.
In this paper, we first investigate and summarize three real-world hurdles in applying known physical-layer localization techniques on UE uplink signals through real-world passive experiments. C1) First, a localizer is able to track a UE only when it generates enough uplink traffic reliably for a certain time period. This means that if a target UE is not actively transmitting signals, the localizer cannot effectively locate its uplink signals. C2) Second, when a target is in close proximity to the cell tower, its uplink signal power may be significantly reduced. This poses a challenge for localizers, as detecting and accurately determining the direction of the target’s uplink signal becomes challenging in urban cellular networks where base stations are densely deployed. C3) Last, cellular repeaters, which are ubiquitous in modern indoor environments, amplify and relay downlink/uplink signals and this introduces severe noise to the localization of true UE uplink signals.
It is noteworthy that cellular operators’ assistance is not sufficient to address the above challenges. Cellular operators have full control over their RAN infrastructure, including base stations (eNBs), repeaters, and UEs; however, their control is limited when it comes to the uplink signals of specific target UEs. RAN vendors offer many management features to mobile operators via proprietary interfaces, allowing them to operate the RAN infrastructure. Nevertheless, these do not extend to making arbitrary changes to the uplink signals of a particular UE. To the best of the authors’ knowledge, no RAN systems exist that allow mobile operators to finely adjust uplink scheduling or power-control, which are necessary to address the challenges of achieving reliable fine-grained cellular localization.
Real-world case: Vishing investigation. As a realistic example, police agents in certain countries trace the location of the cellular devices engaged in vishing fraud called SIM Boxes [22, 23, 44], and subsequently confiscate them [38, 45]. Note that those SIM Boxes are uncooperative devices. They don’t have GPS sensors and only have call-relaying capability. The agents are initially provided with SIM boxes’ coarse-grained location (i.e., cell ID) from mobile operators. They then employ a portable sniffer to track uplink signals emitted by SIM boxes. In this situation, they may encounter the aforementioned challenges, especially in urban cellular networks, and cellular operator’s assistance is insufficient to address them.
To overcome these hurdles, we design an Uncooperative Multiangulation Attack (UMA) and show that overcoming these challenges is indeed feasible with UMA. We show that a localizer with UMA, after obtaining coarse-grained location information of a target UE from the mobile operator, can further obtain UE’s strong uplink signals reliably. We show via end-to-end demonstration in this line of work that real-time manipulation of downlink and uplink signals is feasible and thus physical-layer based uncooperative localization becomes possible with the current, limited operator support. As it requires minimal operator support (e.g., only cell-level information) and works in any LTE networks that follow the mandatory standard features, UMA achieves not only reliable but also universal cellular localization.
UMA involves two main approaches:
(1) We suggest a scheduling manipulation to effectively address C1. With scheduling manipulation, the localizer (e.g. law enforcement), impersonates the target UE and informs the eNB that it has data to transmit. Accordingly, the eNB continuously allocates uplink resources to the target and the target’s connection remains active. Due to constant resource allocation, the target UE continues to transmit uplink traffic even when it has no data to transmit in its current buffer. This scheduling manipulation ensures that the UE is always transmitting, preventing it from becoming idle.
(2) We propose a power boosting, designed to resolve C2 and C3. Using power boosting, the localizer, now, impersonates the eNB and instructs the target UE to increase its transmission power to the maximum level. The power boosting directly addresses C2, as it allows the localizer to increase the target’s signal to its maximum level. Moreover, the power boosting enables the localizer to overcome the impact of a repeater. Since the repeater consistently transmits with constant signal strength, executing the power boosting allows the localizer to observe a notable disparity in signal strength between the target UE and the repeater. Hence, this approach effectively resolves challenge C3 as well.
To sum, UMA has three unique characteristics:
  • Network-wide. Unlike existing techniques, which focus on locating uplink signal sources through physical-layer measurements (e.g., AoA, ToA, TA), UMA addresses yet-overlooked systematic challenges that arise when operating localization techniques without device cooperation.
  • Universal. While existing native 3GPP positioning (e.g., OTDOA via PRS, E-CID) depends on optional features in UE implementations [3], UMA operates effectively in virtually any LTE networks that adhere to the core, mandatory standard features.
  • End-to-end. UMA provides publicly available end-to-end demonstration (e.g., from a target’s phone number to its physical location tracking) of a fine-grained cellular localization technique; see our public demonstration video: https://www.youtube.com/watch?v=Y0MJNQCKs7I.
Note that localization techniques are often regarded as privacy attacks as they could violate user privacy depending on the context in which the techniques are used. While we focus on scenarios that assume authorities with legitimate purposes as the main performer of the localization techniques, we acknowledge that unprivileged third parties may use the UMA technique for unlawful purposes. To simplify writing and our arguments, in this paper, we decide to refer the UMA technique as attacks.
Responsible disclosure. We have reported our findings to GSMA. They have acknowledged our contributions with CVD–2023-70 and CVD–2023-77.
2 Preliminaries
In this section, we provide the technical backgrounds of LTE necessary to understand uncooperative localization.
Identifiers in LTE. When a UE communicates with LTE networks, several identifiers are involved, which are assigned by different network entities. The identifiers can be classified as permanent and temporary. Representatively, the International Mobile Subscriber Identity (IMSI) is a permanent identifier that remains unchanged once assigned to the user’s User Identity Module (USIM).
During radio communication, two temporary identifiers – Globally Unique Temporary Identifier (GUTI) and Radio Network Temporary Identifier (RNTI) – are mainly used to avoid privacy issues [20, 29] caused by the permanent identifiers. GUTI is assigned by core networks when a UE accesses LTE cellular networks and a connection is established. RNTI is assigned by an eNB for physical layer communication in cellular radio networks when the UE requests a connection to radio access networks [4].
RNTI changes periodically whenever a radio connection is re-established between a UE and an eNB. For example, when there is no data transmission between the UE and the eNB for a specific period, the eNB removes the connection for network resource and UE’s battery management. Then, when the UE needs data transmission, a new RNTI is allocated to the UE.
Figure 1: Physical layer channel structure.
Downlink Control Information (DCI). As shown in Figure 1, the eNB and UE transmit their traffic through downlink and uplink channels in the physical layer, respectively. In the physical layer channel, the subframe is a fundamental resource unit for radio communication.
In order to decode downlink data and enable the transmission of uplink data, Downlink Control Information (DCI) plays a crucial role, especially in notifying scheduling information and transmission power control for the UEs. Each UE utilizes the RNTI to decode the DCI message and identify the allocated resources. Specifically, a certain format of DCI, DCI format 0 (DCI 0), is used for controlling uplink transmission power to ensure optimal signal quality and identifying allocated resources (i.e., uplink grant) on Physical Uplink Shared Channel (PUSCH). Note that DCI 0 is broadcast without encryption, it is possible to obtain a UE’s uplink scheduling information by using the UE’s RNTI. The process of DCI decoding and resource allocation is illustrated in Figure 1.
3 Attack Model
3.1 Motivation
Beyond-cell-level, fine-grained localization of cellular devices has been a highly desired feature for various applications, such as emergency response, criminal investigation, and even unlawful stalking or surveillance. Localizers begin by obtaining the coarse-grained location of a target device (e.g., directly obtaining its serving cell ID from the mobile operator [18] or inferring the serving cell via paging side channel [46]), and then attempt to further narrow down.
However, existing studies show that fine-grained cellular localization is hard even for authorized agents with assistance from mobile operators. A survey from Electronic Frontier Foundation in 2021 [21] shows that mobile signal tracking via triangulation by operators offer only about 1 km accuracy. It also discusses other techniques, such as cell site simulators, which are used by law enforcement agencies; however, they allow only coarse-grained device presence detection, and works only when the target device supports 2G, a soon-to-be-obsolete technology in the market. Peral-Rosado et al[17] also have surveyed localization technologies that can be used by mobile operators. It discusses a few technologies that have potential to provide beyond-cell-level localization (e.g., 3GPP’s E-CID, OTDOA, and UTDOA) but they are only optional features [3] and thus cannot provide universal fine-grained localization across all mobile operators, RAN vendors, and UE manufacturers. As an aside, GPS information can be a useful means for authorities to trace the location of the target UE. However, authorities cannot access GPS location without the permission of mobile users when they are using iPhone [36] and devices employed by criminals may lack support for GPS features.
Figure 2: Idealized view vs. real-world view.
Idealized view on cellular localization. Understanding the limited cellular localization capability in modern cellular networks, there have been several academic studies that attempt to achieve fine-grained localization of cellular devices outside the official localization features provided by mobile operators [31, 32]. These studies, however, have rather idealized views on the cellular localization problem, thus overlooking several practical challenges that a localizer may suffer when attempting physical localization in real-world environments.  Figure 2 illustrates this idealized view on cellular localization used in previous studies. In this view, a target device is always transmitting uplink signals, which are always strong enough to be detected by the localizer. Also, strictly only one uplink signal is transmitted by the target device at a time. Thus, the localizer can conveniently conduct physical-layer localization techniques (e.g., AoA, ToA, TA) to locate the target’s uplink signal source, after they obtain the target device’s temporary network identifier (e.g., RNTI) and detect its uplink signal.
While this simplifed view is a good starting point to understand the cellular localization problem, in a more realistic, network-wide view on cellular localization, the localizer immediately faces several challenges because the real RAN environment is much more complex. First, targeted UEs may not be transmitting uplink signals at the time the localizer is attempting to trace them. This not only makes localization hard but also makes it difficult to identify the target UE’s temporary network identifier. Second, target UE’s uplink signal power varies significantly because it is always adjusted depending on its channel condition. Last, the localizer sees not only the target UE’s uplink signal but also many replicas of the same signal amplified and relayed by repeaters that are widely deployed to enhance cellular channel quality.
3.2 Threat Model
We consider an unprivileged attacker, who cannot control the operation of the network infrastructure (e.g., eNB or core networks) or run its own code on target devices. This scenario applies both to authorities (such as police), and attackers. In particular, this setting does also not rely on optional features that are not widely supported in UEs and infrastructure (e.g., 3GPP positioning features).
We assume that the attacker knows an online identity (e.g., phone number) of the victim and the cell information which the victim is camping on. The cell information can be obtained in advance by exploiting the existing presence test [9, 10, 12, 25, 33, 35, 46]. Alternatively, the police agency can request cell information from mobile carriers for their investigation.
Instead of accessing the network and UE directly, the attacker employs multiple devices equipped with directional antennas (i.e., AoA) to measure the direction of the target UE’s uplink signal. Using these devices, the attacker aims to determine the physical location (e.g., 2D-coordinated) of the victim’s UE by measuring the direction of uplink signals at multiple points. In addition, these devices have the capability to passively monitor both uplink and downlink LTE signals. This capability can be achieved using an open-source LTE eavesdropper, such as LTESniffer [24], with off-the-shelf software-defined radio devices (e.g., USRP [51]). One unique capability of the devices is the ability to actively inject malicious LTE signals (i.e., signal overshadowing [19, 54]) on both downlink and uplink channels. Upon obtaining the victim’s uplink scheduling information using these devices, the attacker can employ a directional antenna to measure the direction of the uplink signal source, detecting the strongest signal power. Utilizing Angle of Arrival (AoA) measurements from multiple locations, the attacker performs a multiangulation process to ascertain the precise coordinates of the victim. Alternatively, the attacker can progressively refine the victim’s location by continuously tracking the direction of the uplink signal source (i.e., iterative direction finding).
TABLE I: Uplink transmission power depending on the distance between the eNB and the UE.
Distance (m) 10 20 30 40 50 60 70 80 90 100 110
Tx PWR (dBm) -7.04 -10.64 0.49 4.74 5.05 4.33 7.65 7.45 7.2 7.39 7.56
RSRP (dBm) -62.37 -62.63 -73.43 -76.66 -78.91 -81.69 -85.36 -85.51 -86.45 -84.23 -87.6
Additionally, among various cellular localization techniques, we use AoA measurements of uplink signals as a fundamental technique for physical localization. Other techniques such as TA, ToA, and TDoA have some limitations in employing fine-grained localization of uncooperative UEs in real-world environments. ToA-based cellular localization [31] requires an additional process for fingerprinting the victim’s device to account for hardware errors. This is needed to accurately calculate the arrival signal timing to accurately estimate the victim’s location. In addition, TA-based localization [48] shows only sub-cell-level localization performance with 78m granularity.
3.3 Challenges of Uncooperative Localization
Uncooperative localization attack in real-world environments needs to address the following three challenges:
  • C1) The attacker can track the victim only when the victim’s uplink traffic is generated
  • C2) The attacker is unable to trace the victim when the victim’s signal is undetectable
  • C3) The presence of a cellular repeater can interfere with the accurate determination of the victim’s direction
To confirm the existence of such challenges in operational cellular networks, their impact will be explored using an LTE channel sniffer (Appendix § B.1).
C1: High dependency on victim’s uplink signal. Recall that the primary property exploited by the localization attacker is the monitoring of the uplink signal generated by the victim’s UE. From this perspective, the attacker can initiate localization only when the victim generates uplink traffic. In other words, the localization attack is challenging when the victim is silent (i.e., not transmitting any signal). In such instances, the attacker is compelled to passively wait for the victim’s uplink traffic using their sniffer, as they lack access to either the eNB or the victim UE.
Additionally, it is important to note that RNTI is a temporary identifier that changes frequently (approximately every 15 to 30 seconds [8, 30, 43]). This implies that when a UE is not actively utilizing any services for a specific duration after connection, its radio connection is terminated by the eNB, and the previously assigned RNTI expires. If the UE later requests LTE services from the eNB, a new RNTI value will be assigned for radio communication. The attacker may face a challenge in identifying the victim’s location if their radio connection and assigned RNTI expire during the localization process. This necessitates the attacker to patiently wait for the victim to establish a new radio connection, after which they must track the victim’s new RNTI to resume the localization process.
C2: Undetectable victim’s signal. The successful measurement of arrival signals at the attacker’s sniffer is crucial for executing the localization attack. If the sniffer fails to detect the victim’s uplink signal, estimating the victim’s location becomes challenging. This issue arises because the UE does not consistently transmit its signal with high power to the eNB. Notably, when the UE is situated in close proximity to the eNB, weak uplink transmission power is used. The UE dynamically adjusts its uplink transmission power to compensate for signal path loss, influenced by factors such as distance and obstacles, in coordination with the eNB [2, 37]. For instance, Table I demonstrates how the UE adjusts its transmission power based on both distance and Reference Signals Received Power (RSRP).
The transmission power control between the eNB and UE creates a shadow area where the victim’s signal remains undetectable by the sniffer. To illustrate this shadow area in real-world environments, we examine how the UE’s position influences the sniffer’s signal detection. In our experiment, the attacker’s sniffer, positioned 80m away from the victim, successfully detected uplink signals from victims located 90m away from the eNB. However, when the victim’s UE was only 10m away from the eNB, the sniffer failed to detect uplink signals at the same position (refer to Appendix § B.3 for detailed experimental setup). Unfortunately, due to the attacker’s lack of permission to control networks and UE, they are unable to legitimately instruct the UE to increase its transmission power when encountering the shadow area.
C3: Differentiating the victim from the repeater.
Figure 3: Impact of cellular repeater in localization attack.
A cellular repeater is a device designed to amplify both downlink and uplink cellular signals, reducing signal path loss between the eNB and UE (i.e., improving cellular channel quality). As a consequence of the repeater’s operation minimizing signal path loss, the UE tends to transmit uplink data with lower signal power. Simultaneously, the repeater transmits (i.e., relays) the data with amplified, higher signal power. It is widely deployed to enhance the service coverage of cellular carriers. However, repeaters pose a challenge for attackers in determining the direction of the strongest signal.
Figure 4: Overview of UMA.
[…]
Hence, the attacker needs the capability to discern the direction leading to the UE’s location. Despite its substantial impact on localization performance, none of the previous studies have considered the influence of the repeater.
3.4 Approach: UMA
We present an uncooperative multiangulation attack (UMA), which resolves the key challenges of uncooperative localization (§ 3.3). UMA leverages AoA measurements with directional antennas, composed of six steps (Figure 4). These steps include RNTI acquisition using the victim’s online identity (①–②), manipulation of uplink scheduling (③), boosting the target’s uplink signal strength (④), and measurement of the AoA of the uplink signal (⑤–⑥).
In particular, we come up with two approaches to address the challenges of uncooperative localization; the scheduling manipulation attack and the power boosting attack. The scheduling manipulation attack is specifically designed to address challenge C1, allowing the attacker to acquire the unexpired victim’s RNTI and ensure continuous uplink transmission. To overcome challenges C2 and C3, we employ the power boosting attack, enabling the attacker to increase the victim’s uplink transmission power and distinguish the victim’s signal from the repeater’s. Both attacks exploit vulnerable protocols in LTE specification, which lack implemented security protection measures. By utilizing these techniques, UMA effectively addresses the key challenges associated with uncooperative localization.
Overall procedures of UMA. An UMA attacker follows a systematic process illustrated in  Figure 4. The first two steps are to obtain the victim’s RNTI associated with their phone number (§ 4.1). This is accomplished by sending silent Short Message Service (SMS) messages (①) and monitoring traffic patterns on the downlink channel (②). Note that this RNTI is a key identifier to figure out the victim’s uplink scheduling. Next, the attacker proceeds with a scheduling manipulation attack (③); it transmits fake scheduling requests, and reports fake transmission buffer status to the eNB, posing as the victim UE with its RNTI (§ 4.2). This attack ensures the victim’s radio connection remains active, and their RNTI stays unexpired. Through the scheduling manipulation attack, the attacker forces the victim UE to generate uplink traffic even when there is no data available for transmission in its buffer. Subsequently, the attacker conducts a power boosting attack (④) by injecting manipulated commands for transmission power control logic into the victim’s UE, disguising as the eNB. This attack forces the victim UE to increase its uplink transmission power up to a maximum level (23dBm), overcoming practical challenges caused by the undetectable victim’s uplink signal and the presence of a cellular repeater in the radio network (§ 4.3). After these active operations, the attacker obtains the victim’s scheduling information broadcasted over DCI 0 by monitoring the downlink channel (⑤). With this information, the attacker identifies the victim’s uplink signals on the uplink channel. Then, the attacker measures the AoA of the victim’s uplink signal by rotating a directional antenna mounted on the sniffer (⑥). Ultimately, the attacker can achieve physical localization by tracking down the victim’s physical location iteratively or employing multiangulation with multiple sniffers.
4 Design of UMA
In this section, we present the design of our approaches in a sequential manner. Firstly, we introduce a simple yet feasible method, inspired by previous works, for acquiring the victim’s RNTI using only their online identity (§ 4.1). Next, we describe how an unprivileged attacker compels the victim to generate uplink traffic and maintain the victim’s RNTI (§ 4.2). Finally, we illustrate how the attacker compels the victim to increase their uplink transmission power and overcome cellular repeaters (§ 4.3).
4.1 From Online Identity to Radio Identifier
For an attacker armed solely with the target’s online identity (e.g. phone number, WhatsApp ID), the initial step in UMA is to acquire the target’s temporary identifier used in the radio network (RNTI). The eNB broadcasts uplink scheduling information to each UE, designating a recipient UE by using the RNTI. Consequently, an unprivileged attacker seeking to monitor and ascertain the uplink signal of the victim UE must be aware of the victim UE’s RNTI. Since the RNTI is intentionally designed to be decoupled from the online identity, acquiring the RNTI becomes essential to establish the link between the victim’s online identity and its corresponding RNTI.
Inspired by prior works [25, 27, 30, 33, 35, 43, 46], our approach to acquiring the RNTI involves correlating voice/SMS traffic pattern directed toward the victim with monitored downlink packets identified by the recipient’s RNTI. To this end, we generate successive silent voice calls/SMSes. Also, we leverage our observation that traffic related to control messages generated by incoming voice calls and SMS is distinguishable from user-plane data traffic.
RNTI acquisition is conducted through the following three steps. ➀ Given the victim’s online identity, the attacker makes silent voice calls or sends silent SMSes to the victim following a specific traffic pattern (e.g., multiple times with a constant time gap). ➁ The attacker monitors the downlink traffic of SMS and voice calls. Specifically, they monitor the dedicated channels, called bearers, used for sending control messages of voice calls and SMSes. Note that the attacker is able to monitor the targeted traffic pattern clearly as each bearer is designed to deliver a different type of data. For example, the messages for controlling the voice calls are delivered over a bearer named Data Radio Bearer (DRB) 1, and user-plane traffic and SMS control messages are flooded through DRB 2 and Signalling Radio Bearer (SRB), respectively. ➂ The attacker examines radio connections and finds out the connection showing the intended traffic pattern. Finally, the attacker determines the RNTI of that connection as belonging to the victim.
4.2 Manipulating Victim’s Uplink Scheduling
One of the key pillars of UMA lies in its ability to a) compel the victim to generate uplink transmission and b) retain the acquired victim’s radio identifier (RNTI). To achieve this, we introduce a novel approach known as the Scheduling Manipulation Attack. Essentially, the Scheduling Manipulation Attack injects a crafted message into the eNB, prompting the eNB to schedule resources for uplink transmission to the victim UE and triggering the UE to transmit uplink data. This strategic action prevents the UE from entering an idle state, ensuring the continuity of the acquired victim UE’s radio identifier (i.e., RNTI) obtained from the victim’s online identity (§ 4.1).
4.2.1 Uplink Scheduling Procedure
We begin by outlining the control mechanism governing a UE’s uplink data transmission. Essentially, the UE is granted permission to transmit uplink data only when it receives uplink grants allocated by the eNB. The UE follows a defined procedure to secure uplink grants for data transmission to the eNB. Initially, the UE transmits a scheduling request [2], signaling to the eNB its intention to transmit uplink data. Upon receiving the scheduling request, the eNB grants approval for the UE’s Buffer Status Reporting (BSR) [4], a mechanism used to report the size of uplink data. Subsequently, the UE initiates the BSR, leading to the allocation of sufficient uplink grants by the eNB, allowing the UE to transmit its data.
Scheduling Request is a PHY layer message sent from the UE to the eNB over Physical Uplink Control Channel (PUCCH). To transmit a scheduling request to the eNB, the UE uses schedulingRequestConfig [2, 5] previously shared via the RRC connection setup message. schedulingRequestConfig notifies each UE of which resource should be used to transmit scheduling requests.
Buffer Status Reporting (BSR) is a MAC layer message sent over PUSCH to notify the eNB of the data awaiting transmission in the UE’s buffer. Following the reception of the BSR, the eNB issues multiple uplink grants to the UE, ensuring it has ample resources to transmit the queued data.
4.2.2 Design of Scheduling Manipulation Attack
Figure 5: Procedure of scheduling manipulation attack.
Our approach exploits BSR [48] and scheduling requests to induce unintended uplink transmissions and maintain an unexpired radio connection. Scheduling requests and BSR, governed by the PHY and MAC layers, lack security protection such as encryption or integrity checks. The core idea is to manipulate uplink scheduling by falsifying the victim’s uplink channel, prompting the eNB to allocate uplink grants to the victim, even when there is no actual uplink data to transmit. Our approach unfolds through three distinct steps, as illustrated in Figure 5.
1) Acquiring user-dedicated radio configuration. The UMA attacker aims to identify the victim’s user-dedicated radio configuration. The user-dedicated radio configuration, which includes the schedulingRequestConfig, is communicated to each UE by the eNB during the establishment of a radio connection. This information is conveyed through the RRC Connection Setup message. The attacker can discern the victim UE’s configuration by monitoring the RRC Connection Setup when the victim attempts to establish a radio connection with the eNB.
2) Establishing forged uplink channel. The UMA attacker manipulates the victim’s uplink channel by utilizing the acquired radio configuration and RNTI of the victim. Initially, the attacker establishes a radio connection with the eNodeB to create a communication channel. Subsequently, the attacker configures its RNTI, and schedulingRequestConfig [2, 5] with identical values to those of the victim. This setting allows the attacker to gain control of the victim’s communication channel.
3) Manipulating uplink scheduling. After acquiring the victim’s communication channel, UMA attacker injects fake uplink scheduling request and fake BSR to the eNB to manipulate uplink scheduling. First, a fake scheduling request is transmitted over the victim’s resource on PUCCH, which is obtained by forging the victim’s uplink channel. Then, the eNB allocates uplink resources to the victim’s RNTI on PUSCH. Second, to obtain multiple uplink grants, the attacker transmits fake BSR over the allocated resource through fake scheduling request. The fake BSR consists of a) the victim’s RNTI, b) the required buffer size, and c) the Logical Channel ID (LCID). In this attack, we set the required buffer size to 200 bytes. The LCID has a value of 3 which indicates data bearer communication in uplink logical channels. The attacker repeatably conducts the above process to make the eNB continuously allocate uplink grants to the victim. At this time, manipulated scheduling requests and BSR are transmitted only using 10% of allocated uplink resources to minimize the impact on the victim’s Quality of Service (QoS).
4.3 Boosting Victim’s Uplink Signal
In this section, we present a novel approach called Power Boosting Attack to enforce a UE to increase its uplink transmission power. This attack is required to overcome a shadow area where the victim’s signals are undetectable to the localization attacker and the impact of a cellular repeater. We first introduce a mechanism for the eNB to control the UE’s uplink transmission power (§ 4.3.1). We then present the implementation of the power boosting attack exploiting TPC commands (§ 4.3.2).
4.3.1 Transmit Power Control
TABLE II: $\Delta_{PUSCH}$ according to TPC command in DCI 0.
TPC command in DCI 0 0 1 2 3
Accumulated $\Delta_{PUSCH}$ (dB) -1 0 +1 +3
An eNB is responsible for all aspects of data transmission with the connected UEs. Along with the data transmission scheduling, it also controls the data transmission power of the UE by Transmit Power Control (TPC) command, contained in DCI 0 messages. The UE adjusts its transmission power based on the internal power control algorithm, which uses a designated value in the TPC command along with the current self-monitored channel quality [2]. One widely used algorithm adopts the accumulated power control that gradually adjusts the transmission power according to TPC commands as defined in Table II, with a maximum power of 23dBm.
4.3.2 Design of the Power Boosting Attack
Our novel approach to changing the target UE’s uplink transmission power without any privilege is to utilize this TPC command. The TPC commands are delivered over DCI 0 messages without any security protection (i.e., encryption or integrity check). In other words, it is exposed to a downlink overshadowing attack [54]. The key idea is to inject manipulated subframes that contain DCI 0 broadcast to the victim UE, whose TPC is set to increase the victim UE’s transmission power. This power boosting attack can be realized by following three concrete steps.
1) Acquiring configuration for radio communication. To identify the structure of a legitimate downlink subframe, the UMA attacker retrieves the physical configuration of the target eNB. This configuration includes target Physical Cell Identification (PCI), channel bandwidth, Physical channel HybridARQ Indicator Channel (PHICH) configuration, cyclic prefix, and transmission mode of the eNB. This can be retrieved by decoding the messages that the eNB broadcasts such as Master Information Block (MIB), Primary Synchronization Signal (PSS), and Secondary Synchronization Signal (SSS).
2) Manipulating subframe. UMA attacker crafts a subframe that includes a malicious DCI 0 message using the obtained physical configuration. The manipulated DCI 0 messages then consist of a) the victim’s RNTI, b) arbitrary scheduled uplink transmission information, and c) the TPC command having a value of 3 to boost the uplink transmission power.
3) Overshadowing the manipulated subframe. By extending the signal overshadowing attack [54], UMA attacker injects the crafted subframe with the proper subframe number over PDCCH. More precisely, we mainly modify the subframe construction logic of the SigOver attack and use almost the same logic for injecting the resulting subframe. In this work, we inject the manipulated subframe to overshadow every legitimate subframe that has index number 9. This is mainly because 1) we want to minimize the impact on the victim’s QoS, and 2) there is a race condition with the eNB. As the eNB also tries to adjust the UE’s uplink transmission power, the UE decreases its power whenever it receives the DCI 0 messages from the eNB through the remaining subframes. Therefore, we repeatedly conduct the overshadowing attack with the manipulated subframe.
4.3.3 Strategy for defeating cellular repeater
UMA could distinguish the victim’s signal from the repeater’s signal by employing a power boosting attack. The key idea is to exploit the cellular repeater’s operational logic, specifically designed to amplify the input signal with its maximum output power [^1] The maximum transmission power of the repeater is typically specified in its product documentation., enhancing the overall LTE channel quality. Thus, even after subjecting the victim UE to a power boosting attack, the uplink signal amplified by the repeater remains unchanged.
In this context, the attacker conducts the following procedures. First, the attacker injects a manipulated TPC command having a value of 3 over DCI 0 message. Second, both the victim UE and the repeater receive that message. Third, the victim UE increases its uplink signal strength, but the repeater does not, as it already transmits the uplink signal with its highest power. As a result, it leads to an increase in the arrival signal strength only on the victim UE side, as observed from the sniffer’s perspective. Based on this strategy, the UMA attacker can correctly determine the victim UE’s direction, overcoming the practical challenge posed by the presence of the cellular repeater.
This strategy is viable due to the distinct operational logic of the User Equipment (UE) and the repeater. The repeater serves as a passive network component, solely amplifying the input signal with its maximum output power before relaying the amplified signal. In contrast, the UE dynamically adjusts its uplink transmission power based on the signal path loss, a topic discussed in § 3.3. Consequently, when subjected to a power boosting attack, the UE and repeater exhibit different responses. The repeater consistently amplifies the signal with its maximum power output, while the UE adapts its transmission power dynamically. This discrepancy in the reactions to power boosting allows the attacker to distinguish and identify the victim’s signal effectively.
TABLE III: Test environments for each experiment.
Experiment Environment
RNTI acquisition
ON (§ 5.1)
Scheduling manipulation attack
LT (§ 5.2), CT (§ 5.4)
Power boosting attack
LT (§ 5.3), CT (§ 5.4)
End-to-end evaluation of UMA
LT (§ 5.5)
Defeating cellular repeater
ON (§ 5.6)
Iterative direction finding of a UE
ON (§ 5.7)
Multiangulation using uplink signals
ON (§ B.2)
Existence of shadow area
ON (§ B.3)
Received strengths for LTE services
ON (§ B.4)
  • ON: Operational network, LT: Lab testbed using srsRAN, CT: Commercial testbed
TABLE IV: COTS UE used in the experiments.
Device Baseband Lab testbed Commercial testbed
Galaxy Note FE Exynos 8890 S+P -
Galaxy Note 10 5G Exynos 9825 S+P -
Galaxy S10 5G Exynos 9820 S+P S+P
Galaxy S20 Snapdragon 865 S+P S+P
LG G8 Snapdragon 855 S+P -
Huawei P30 $*{}^{*}$ Kirin 980 S -
Galaxy Z Flip4 Snapdragon SM8475 S+P S+P
iPhone $*{}^{*}$ Intel PMB 9955 S -
Redmi Note 9T $*{}^{*}$ Mediatek Dimensity 800U S -
  • S: Scheduling manipulation attack (§ 5.2 and § 5.4), P: Power boosting attack (§ 5.3 and § 5.4)
  • *** Power boosting attack could not be confirmed on these devices due to the incompatibility of their baseband chipsets with our XCAL [53], which is required to monitor uplink transmission power. However, given that their baseband chipsets are also designed to follow 3GPP standards, it is expected that power boosting attack should work on these devices.
5 Evaluation of UMA
We demonstrate the UMA in three different testing environments; a lab testbed (LT), a commercial testbed (CT), and an operating commercial network (ON). Table III summarizes conducted experiments. First, we validate the RNTI acquisition (§ 5.1) in a busy commercial network, affirming the feasibility of scheduling manipulation (§ 5.2) and power boosting (§ 5.3) in the lab testbed. Subsequently, we demonstrate scheduling manipulation and power boosting attacks in the commercial testbed, highlighting their practicality (§ 5.4). An end-to-end evaluation of UMA is then presented, showcasing how the proposed approaches can be integrated to perform the physical localization of uncooperative cellular devices (§ 5.5). Subsequently, we show that power boosting attack can be used to distinguish the UE from the cellular repeater in the commercial network (§ 5.6). Finally, we provide a real-world scenario demonstration where the adversary tracks down the target from outside the building to the front door in an urban commercial network through iterative direction finding (§ 5.7).
Experimental environments. The lab testbed is built using an open-source LTE platform, srsRAN [47], supporting a full-stack LTE eNB. The commercial testbed, established by a national institute, comprises an industry-grade LTE solution from Nokia, situated in a shielded room measuring 3m $\times$ 10m. In addition, we used the operational commercial network of a mobile carrier with the largest subscriber base in [our country]. We evaluated scheduling manipulation and power boosting using nine COTS devices (Table IV) in testbed environments (Table III), due to ethical considerations. Both attacks were implemented by modifying srsRAN [47] with USRP X310 [51], following the aforementioned steps in § 4.2.2 and § 4.3.2.
5.1 RNTI Acquisition
To validate the proposed RNTI acquisition procedure, we assess its ability to identify the victim’s RNTI based on their online identity, even in a crowded operational network.
Initially, we investigate the existence of a distinguishable voice/SMS traffic pattern. To do so, we build a dataset comprising over 4 hours of downlink traffic from two operational LTE networks in two countries. This dataset involves 10–60 active users connected to the eNB and a total of 2,142 radio connections. We could observe that, as an example case, none of the users has a traffic pattern featuring multiple messages/calls with a time gap of over 6 seconds. This observation suggests that the attacker could identify the target victim’s RNTI by carefully designing a transmission strategy for the distinctive traffic pattern.
With the identified traffic pattern and the target’s online identity, we execute the RNTI acquisition process in a crowded commercial network with 862 active users detected over an hour. Using an iPhone 14 Pro as the sender and a Samsung Galaxy S10 5G as the receiver, we conduct this trial ten times. We could confirm that the connection identified by the target’s RNTI is distinguishable from others, achieving a 100% success rate over all trials.
5.2 Scheduling Manipulation Attack
We investigate the effectiveness of manipulating uplink scheduling in UMA. To demonstrate the feasibility of the scheduling manipulation attack, we conduct evaluations using nine COTS devices in the testbed environment (Table IV). To prevent the UE from requesting uplink grants during the attack, we disable all applications, including background processing.
First, we assess whether the victim’s UE generates uplink traffic even when no data is available for transmission, and uplink grants are allocated by the eNB. To inspect the UE’s uplink transmission, we collect uplink packets of the MAC layer at srsENB during scheduling manipulation. Next, we investigate whether the victim’s RNTI remains unchanged during the manipulation of uplink scheduling. Throughout the attack, we monitor the radio status of the UE on both the srsENB and UE sides.
Surprisingly, we verify that the UE transmits uplink traffic to the eNB whenever uplink grants are allocated, even when these grants are assigned through manipulated scheduling requests and BSR (and it has no data to transmit in its buffer). Additionally, through the scheduling manipulation attack, the radio connection between the UE and eNB is consistently maintained and the RNTI of the target UE remains unchanged.
5.3 Power Boosting Attack
We investigate whether the TPC injection indeed increases UE’s transmission power up to the maximum level. To demonstrate the effectiveness of the power boosting attack in UMA, we run the experiment in the testbed environment, using six COTS devices, as described in Table IV.
In this evaluation, we initially execute scheduling manipulation attack before injecting TPC commands to show the feasibility of an integrated approach. Before and after the attack, we measure the transmission power of the UE using a diagnostic monitoring tool, XCAL [53], to show the increase of uplink transmission power.
As a result, we confirm that while UE’s transmission power is about 10dBm before performing the attack, through the power boosting, the transmission power of UE is increased up to 23dBm (i.e., maximum transmission power). Considering the transmission power of the UE presented in Table I, it is feasible to overcome the shadow area by increasing the transmission power to 23dBm through the power boosting attack.
5.4 Demonstration in Commercial Testbed
To demonstrate the practicality of scheduling manipulation and power boosting attacks, we examine the combined attack in the commercial LTE testbed. In this evaluation, we execute scheduling manipulation attack as the first step to compel the UE to generate uplink transmissions and maintain the UE’s RNTI. Subsequently, we conduct power boosting attack to increase the UE’s uplink transmission power while the scheduling manipulation attack is ongoing. In this experiment, we use three COTS devices (Table IV). As a result, we successfully conduct scheduling manipulation and power boosting attacks on all UEs in the commercial testbed. Demonstration videos illustrating the combined attack in commercial testbed are available [^2] Demo videos: S10 – https://www.youtube.com/watch?v=Z_b9eR65Q50 S20 – https://www.youtube.com/watch?v=KC2Baoi_1do ZFlip4 – https://www.youtube.com/watch?v=x02KNdxEosU&t=30s .
5.5 End-to-End Evaluation
Now, we demonstrate an end-to-end evaluation on UMA, where an adversary executes four steps in a sequence: 1) RNTI acquisition, 2) Scheduling manipulation, 3) Power boosting, and 4) Physical localization (i.e., multiangulation).
Experimental Setup.
Figure 6: Experimental setup for end-to-end evaluation.
We conduct the experiment in $15m\times 10m$ lab environment as shown in Figure 6. Note that, due to ethical concerns regarding scheduling manipulation and power boosting attacks, we built the testbed with srsRAN [47]. Even though we install srsENB using high-performance SDR, USRP X310 [51], its cellular range is inevitably limited. Therefore, we arrange the experimental setup in a spatially constrained room. We deploy two sniffers equipped with the directional antenna [28], one near the eNB and another at the border of the cellular range, respectively. As depicted in the figure, the target UE is at nine points. For each evaluation trial, we turn on extra two UEs at the same time, connecting them to the eNB. Afterward, our evaluation is designed to figure out where one of the UEs (i.e., target UE) is physically located. Especially, since srsENB doesn’t support SMS, we utilize WhatsApp for RNTI acquisition. More specifically, in this step, we generate traffic patterns without triggering any alarms on the target UE using the ‘typing notification’ feature of WhatsApp [46].
Results.
TABLE V: Evaluation of end-to-end experiment.
Localization performance
Evaluation metrics w/o power boosting attack w/ power boosting attack
Success rate 77% 100%
70% angluar accuracy $\sim$ ${}^{\circ}$ $\sim$ ${}^{\circ}$
70% distance accuracy $\sim$ 2.8m $\sim$ 1.7m
Arrival signal strength (Max.) -2 $\sim$ 19 dBm 33 $\sim$ 38 dBm
Time consumption for each step
RNTI acquisition $\leq$ 30 seconds
Scheduling manipulation attack $\leq$ 30 seconds
Power boosting attack $\leq$ 2 minutes
Multiangulation $\leq$ 2 minutes
Table V shows the summarized performance results from the end-to-end evaluation of the various metrics. We confirm that all nine trials of end-to-end evaluation successfully pinpoint the physical location of the target UE with 70% distance error of 1.7m. Furthermore, we confirm that each UMA attack, where the four key steps are conducted in sequence, takes less than 5 minutes to determine the physical location of a single target UE from its online identity. Not surprisingly, it takes about 2 minutes to perform power boosting attack due to the process of acquiring radio configuration and manipulating subframes, as described in § 4.3.2. In multiangulation step, it also takes about 2 minutes to estimate the direction (i.e., angle) of the signal source at two sniffers and determine the location.
We further analyze the effectiveness of the power boosting attack on the end-to-end evaluation. Table V presents a performance comparison between two scenarios: the end-to-end evaluation with and without the power boosting attack. The power boosting attack enhances localization performance, as evidenced by improvements in angular accuracy, localization accuracy, and arrival signal strength on the sniffers’ side. Additionally, in cases where we perform the end-to-end process without executing the power boosting attack, the physical location of the target remains undetermined at two out of nine locations, marked by gray circles in Figure 6.
To investigate these differences, we measure the arrival signal strength according to the angle from the UE placed at the left gray circle in Figure 6. We rotate the antenna from ${}^{\circ}$ (i.e., pointing to the UE) to ${}^{\circ}$, using the sniffer placed at the bottom right. As depicted in Figure 7, in the absence of the power boosting attack step, the angle corresponding to the peak signal strength is not discernible, making it challenging to determine the angle of the signal source. This directly leads to errors in multiangulation, causing the attacker to pinpoint the physical location inaccurately. This observation aligns with the discussion in § 3.1, where it is noted that when the UE is in close proximity to the eNB, it generates its uplink signal with low transmission power. Additionally, Figure 7 hints that power boosting can improve the direction finding capability by reducing the noise on the sniffer side when the attacker attempts to estimate the direction of signal source. Actually, the signal-to-noise ratio (SNR) in the UE direction has been found to be 30dB with power boosting and 13dB without power boosting. A demonstration video illustrating the end-to-end evaluation is available: https://www.youtube.com/watch?v=Y0MJNQCKs7I&t=1s.
Figure 7: Impact of the power boosting on direction finding.
5.6 Power Boosting vs. Cellular Repeater
We now examine the effectiveness of our strategy employing the power boosting attack to distinguish the UE’s signal from the repeater’s. Due to ethical concerns, the evaluation is conducted on the commercial testbed. This testbed comprises commercial network equipment (§ 5). The repeater’s external antenna was installed near eNB and the UE was positioned 2m away from the repeater. We then placed the sniffer 5m away from the UE and the repeater’s external antenna.
We first measure the arrival signal strength at the direction of the UE and the repeater’s external antenna (Table VI) without the power boosting attack. The arrival signal strength at the direction of the UE is measured as 5dBm. At the direction of the repeater, the signal strength is equivalent to 17dBm. Next, we measure the arrival signal strength in two directions, as before, while conducting the power boosting attack. Signal strength at the direction of UE, at this time, significantly increases to 25dBm, while the arrival signal strength at the direction of the repeater remains unchanged (static). As a result, as we described in § 4.3.3, by observing an increase in arrival signal strength using a directional antenna after performing the power boosting attack, the UMA attacker can identify the UE’s true signal.
TABLE VI: Impact of cellular repeater and TPC injection.
TPC injection Off On
Target UE Repeater UE Repeater
Signal strength (dBm) 5 17 25 17
  • Distance between the repeater’s antenna and the UE is 2m
In conclusion, through the observation of an increase in arrival signal strength using a directional antenna after performing the power boosting attack, the UMA attacker can distinguish the location of the UE from that of the repeater.
5.7 Iterative Direction Finding in operational network
Lastly, we present a use case for the physical localization of uncooperative cellular devices in an operational LTE network. More specifically, similar to the scenario we described in § 3.1, we show how a localizer can find a UE in a room on the 2nd floor of a building measuring $56m\times 34m\times 15m$, camped in the urban operational network. The demostration begins by deploying a sniffer near the eNB (approximately 10m away from the tower).
To acquire the target’s RNTI, we employ SMS messages to generate traffic patterns. This is achievable by monitoring downlink traffic, as discussed in § 4.1. After acquiring the target’s RNTI, scheduling manipulation would be conducted to force the UE to generate uplink traffic and maintain the UE’s RNTI. Due to ethical considerations, however, we employ voice call service to replace the effect of scheduling manipulation and power boosting. Naturally, a voice call makes UE continuously generate uplink traffic and maintain its radio connection (i.e., fixed RTNI). Furthermore, UE’s transmission power increases in response to data throughput requirements of voice call – we empirically confirm this in § B.4 along with the specification [1]. Subsequently, we track down the physical location of the target UE by moving the sniffer’s position.
Figure 8: The process of iterative DFing and localization heatmap targeting UE inside a building.
Figure 8 illustrates the tracking process in this scenario. Initially, we estimate the direction of the signal source at the starting point (L1) and deduce the building where the target UE is suspected to be located. Then, we move the sniffer to the front of the suspected building (L2). At the second point, we investigate on which floor the target UE is located by measuring arrival signal strength, as illustrated at the bottom of Figure 8. Based on the measurement, we determine that UE is most likely on the 2nd floor. Finally, after moving the sniffer to the 2nd floor inside the building (L3), we successfully pinpoint the physical location of the target UE by tracking down the direction of signal source. During the tracking process, if the target UE’s uplink signal is undetectable or the repeater is installed, power boosting can be performed to increase the target’s transmission power.
6 Related Work
TABLE VII: Comparison with related works
ToA-based approach AoA-based approach
RNTI tracking from online identity [31]${}^{1}$ UMA
Targeted localization [31, 48] UMA
No need for H/W or RF fingerprinting [48]${}^{2}$ UMA
Generating continuous uplink signal UMA
Boosting uplink signal power UMA
Consider cellular repeater UMA
[…]
  • 1
    LTrack obtains RNTI from IMSI/TMSI, which are more challenging to acquire compared to online identities.
  • 2
    Localization performance with 78m granularity.
  • 3
    They utilize both passive and active approaches for localization.
Several studies proposed the physical localization methods against cellular devices [31, 32, 48]. These prior works and our work utilize the downlink/uplink signal for localizing the target UE. However, our work has additional capabilities in its threat model and provides novel approaches to address the challenges that have not been discussed before (refer to Table VII). Note that previous works have different goals and threat models.
ToA-based localization. Kotuliak et al[31] proposed a localization attack based on ToA measurement. Their approach exploits the unencrypted information for the transmission time alignment (i.e., TA) and the actual monitored uplink transmission time. Based on these two values, the attacker estimates the victim’s location with a 90% probability and an error of less than 6m. Tan et al[48] proposed data-plane signaling forgeries and utilized the fabricated malicious signaling to determine the position of IoT devices. Similar to the above work [31], they estimated the target’s location based on the eavesdropped TA values from multiple eNBs. However, their method, which solely relies on TA commands, shows insufficient precision (78m).
Compared to those works, our work differs in three aspects: (1) our scenario leverages the victim’s online identity, which is more easily obtainable compared to IMSI/TMSI, for RNTI tracking; (2) we introduce a scheduling manipulation attack to enforce the victim UE to transmit an uplink signal, ensuring a fixed RNTI. This approach helps overcome practical challenges during the localization attack; (3) the challenges and evaluations are built upon a real world scenario. This is important as those challenges are only identifiable in the practical environment.
AoA-based localization. The closest work to our work in the context of localization method (i.e., AoA) is LTEye [32]. The authors introduced an indoor localization attack based on AoA measurement, which utilizes a synthetic aperture radar to handle the multipath problems in indoor environment. They estimated the target UE’s location by using multiple radars at different locations. LTEye also provided an RNTI tracking method to map RNTIs assigned to the same device, which correlates the currently monitored RNTIs to the expired RNTI. However, this method necessitates an RF fingerprinting procedure with a corresponding database. While both LTEye and UMA rely on uplink signal monitoring, UMA employs a more active attacker model by exploiting the uplink transmission logic through the power boosting attack and scheduling manipulation attack. In contrast, LTEeye solely depends on passive monitoring of uplink traffic. Note that, the two proposed novel attacks enable more stable location tracking for both LTEye and UMA. In addition, UMA is designed to targeted attacks, but LTEye cannot be utilized in targeted localization attacks (i.e., estimating the location of a specific victim) since it cannot identify a specific device.
Other approaches. Lakshmanan et al[34] demonstrated unprivileged location tracking by exploiting Carrier Aggregation (CA) side channels in cellular networks. However, it requires collecting CA measurements in all possible target paths, limiting the capability of localizers when attempting to trace cellular devices. Besides cellular localization, Abedi et al[7] presented uncooperative WiFi localization through the injection of fake packets. Another example, Blue’s Clues [50] introduced a method to track Bluetooth devices even when they are in non-discoverable mode. In addition, Zhang et al[57] presented the physical localization of cellular devices by impersonating the victim and replaying messages to the victim’s Bluetooth device. Nonetheless, since these methods exploit short-range communication protocols, the attacker’s ability to trace the victim is limited to a constrained range.
7 Discussion
Novelty and goal. Our main contribution is identifying and addressing realistic challenges associated with physically tracking uncooperative devices in cellular networks through two novel approaches; scheduling manipulation and power boosting. While various localization techniques for precise tracking have been actively studied, our perspective on achieving reliable uncooperative cellular localization has not been previously discussed. We have demonstrated the feasibility of our approaches through evaluations in three environments, within legal boundaries. Based on our findings and evaluations, GSMA has acknowledged our contributions and confirmed that they are exploitable for localization.
While we are not primarily focusing on introducing new localization techniques or enhancing performance, our approach presents an opportunity for performance improvement. In this work, although we demonstrated the effectiveness of UMA against the received signal strength-based localization, our approach is also applicable to Channel State Information (CSI) based localization methods [40, 56, 16], as well as other physical-layer techniques [31, 32]. In such integrated scenarios, the performance of localization may be enhanced by mitigating noise and interference through scheduling manipulation and power boosting, as more signals with higher signal gains tend to result in better channel quality.
Although BSR exploitation, a key component of the scheduling manipulation attack, has been discussed in [48], our approach differs in terms of attack procedures and goals. In [48], the attacker attempts to achieve uplink resource draining by sending forged BSR repeatedly, as a single-protocol attack. In our approach, the attacker impersonates the victim UE and maliciously injects both scheduling requests and BSR of the victim to achieve two objectives necessary for uncooperative physical localization: 1) compel the victim to continuously generate uplink traffic and 2) maintain the victim’s RNTI.
UMA with prior studies. LTEye [32] is a representative study employing AoA for cellular localization, similar to UMA. Thus, our approaches are easily applicable to LTEye. As an initial step of targeted localization with UMA and LTEye, the attacker acquires the target UE’s RNTI associated with its online identity by generating a silent traffic pattern. Then, the attacker executes the scheduling manipulation attack to force the UE to continuously generate uplink traffic and maintain its RNTI. This enables LTEye to effectively perform uncooperative physical localization without its practical requirements; 1) the target UE should be in an active status while the SAR’s mobile antenna rotates 360°and 2) the RNTI mapping process with RF fingerprinting is needed when the target’s RNTI is reassigned. When the attacker faces challenges stemming from low uplink transmission power or the presence of the repeater, the attacker performs the power boosting attack to improve localization performance.
While we utilized multiangulation-based localization, the proposed approaches can also benefit multilateration-based localization. For instance, we can improve LTrack [31] utilizing ToA-based multilateration, by addressing challenges such as tracking RNTI associated with online identity instead of IMSI extraction, ensuring continuous uplink transmission during multiple ToA measurements, and achieving ToA in the case with low signal power. Furthermore, attackers employing LTrack can overcome the impact of repeaters by differentiating the signal through power boosting.
Thanks to UMA, they can achieve reliable fine-grained localization in realistic scenarios such as vishing investigation.
Multipath effects. The proposed attack considers realistic scenarios, yet additional challenges might arise in multipath environments due to frequent reflections caused by obstructions. Fortunately, existing research has explored strategies for radio direction detection [49, 52] and indoor localization [14, 39, 55, 32]. Thus, we believe integrating those algorithms can make localization accurate in such environments.
Countermeasure. First, one plausible method to mitigate a power-boosting attack is to encrypt the radio traffic carrying DCI 0. Second, as a potential solution to prevent scheduling manipulation attack, allocating the radio configuration for scheduling request over an encrypted channel can be considered. One possible approach is to reallocate user-dedicated radio configuration, using RRC Connection Reconfiguration. By doing so, it becomes challenging for the attacker to identify the victim’s schedulingRequestConfig, which is crucial for manipulating uplink scheduling. While UMA is a permanent and unpatchable attack due to the finalization of the LTE specification, ongoing 5G/6G projects can potentially mitigate the threat posed by UMA with these countermeasures.
8 Conclusion
In this paper, we present UMA, enabling the physical localization of uncooperative cellular devices. Our work shows that a localizer with only the target’s online identity can determine the physical location of the target by actively exploiting vulnerabilities in LTE network protocols. Through ensuring strict compliance with legal boundaries, we have demonstrated the attack feasibility and effectiveness of UMA. We emphasize that UMA is a permanent and unpatchable attack since it utilizes vulnerabilities of LTE networks of which specification development is complete.
Appendix A Acronyms
3GPP Third Generation Partnership Project
AoA Angle of Arrival
BSR Buffer Status Report
CDF Cumulative Density Function
CID Cell ID
COTS Commercial off-the-shelf
CVD Coordinated Vulnerability Disclosure
C-RNTI Cell RNTI
DCI Downlink Control Information
DFing Direction Finding
DL DownLink
DRB Data Radio Bearer
eNB Evolved Node B
E-CID Enhanced Cell ID
GSMA Global System for Mobile Communications Association
GUTI Globally Unique Temporary Identity
LCID Logical Channel ID
LTE Long Term Evolution
MAC Medium Access Control
MIB Master Information Block
MUSIC Multiple Signal Classification
OTDoA Observed Time Difference of Arrival
PDCCH Physical Downlink Control CHannel
PDSCH Physical Downlink Shared Channel
PHICH Physical HybridARQ Indicator Channel
PHY Physical
PSS Primary Synchronization Signal
PUCCH Physical Uplink Control Channel
PUSCH Physical Uplink Shared Channel
QoS Quality of Service
RF Radio Frequency
RNTI Radio Network Temporary Identifier
RRC Radio Resource Control
SDR Software Defined Radio
SMS Short Message Service
SRB Signaling Radio Bearer
SSS Secondary Synchronization Signal
TA Timing Advance
TDoA Time Difference of Arrival
TMSI Temporary Mobile Subscriber Identity
ToA Time of Arrival
TPC Transmit Power Control
Tx PWR Transmission Power
UE User Equipment
UL UpLink
UMA Uncooperative Multiangulation Attack
UTDoA Uplink Time Difference of Arrival
Appendix B Localization Attack in the Wild
B.1 LTE Channel Sniffer
Figure 9: LTE channel sniffer.
To get hands-on experience with the localization attack with multiangulation, we first build the sniffer. Then, to examine the performance of the attack in the wild, we conduct localization in the commercial cellular network.
Note that we do not use specialized hardware such as LTEye [32]. Instead, we use a parabolic antenna [11] to eavesdrop on RF signals of downlink and uplink channels. The resulting LTE channel sniffer runs on an Intel Core i5 desktop with SDR. As shown in  Figure 9, the left channel of the USRP is used for the downlink channel eavesdropping, and the right channel is for the uplink channel eavesdropping. A directional antenna is mounted on the uplink channel of the USRP, and both channels share the same clock.
B.2 Experiment of Localization Attack
Figure 10: Estimated location of the UE using three sniffers.
(a) $\theta$ Accuracy
(b) Distance Accuracy
Figure 11: Angular and distance error of estimated location.
Using the LTE channel sniffer, we conduct localization with three sniffers and use a COTS device (i.e., Galaxy S10) as a target. The sniffers and the UE are deployed in the same cellular coverage. The experiment is performed in a Line-of-Sight (LoS) environment of $100m\times 100m$ size and we deploy three sniffers, as denoted by the purple square in Figure 10. The UE is deployed at 30m intervals in nine places, shown as red circles in Figure 10.
First, we estimate the direction of the strongest uplink signal detected at each sniffer while rotating the directional antenna (i.e., AoA measurement). Then, we perform multiangulation based on AoA measurements of each sniffer to predict the location of the UE. The predictions for the UE’s location at each point are denoted as three yellow crosses in Figure 10. Since we already know the location of the sniffers, we could estimate the location of the target UE in coordinate units.
We evaluate the performance of AoA measurements with a cumulative density function (CDF) of angular error as shown in Figure 11 (a). The maximum angular error of AoA is about ${}^{\circ}$ and the minimum angular error is about ${}^{\circ}$. This might include human error during the process of rotating the antenna and estimating the direction of the strongest uplink signal. Then, we evaluate the performance of location predictions with a CDF of distance error. As shown in Figure 11 (b), the maximum distance error is about 14m and the minimum distance error is about 10m. When two sniffers are used to estimate the location of an UE, it shows better localization performance than the case of combining the measurements of three sniffers.
B.3 Shadow Area: Impact of the UE’s Position
Figure 12: Arrival signal strength according to the distance between the UE and the sniffer.
We measure the arrival signal strength at the sniffer, which runs on a 10Mhz bandwidth LTE channel in an 800Mhz band. We use a COTS device, Samsung Galaxy S10, as the victim’s UE. Both the UE and the sniffer are connected to the commercial network and located within the same commercial cell coverage in an urban environment. In this experiment, we set the parabolic antenna [11] mounted on the sniffer to direct to the UE. To calculate the signal strength, we take an average of the monitored strength at the sniffer over 2,000 subframes during each measurement trial.
First, we consider the case in which the UE is located far from the eNB. We place the UE 90m away from the eNB. We then place the sniffer in the middle of the eNB and the UE. The sniffer measures the arrival signal strength every 10m between the eNB and the UE. Second, we consider the case when the UE is located close to the eNB, in contrast to the first case. For this, we place the UE 10m away from the eNB. The sniffer is situated 20m from the eNB, which is 10m from the UE. Then, we measure the signal strength at the sniffer by placing it every 10m up to 80m from the UE.
In Figure 12, orange-colored and blue-colored lines show the measured signal strength at the sniffer. When the UE is far from the eNB (orange line), the attacker’s sniffer is able to detect the victim’s uplink signals well, up to 80m away from the UE. However, when the UE is close to the eNB (blue line), the sniffer can no longer detect the victim’s uplink signals in the case the sniffer is situated more than 70m away from the UE. This is because the UE transmits its uplink data with lower signal power when the UE is located close to eNB.
B.4 Impact of the UE’s Behavior
Figure 13: Arrival signal strength by the UE’s behavior.
We investigate the impact of a UE’s behavior on the attacker’s sniffer, the UE using the three behaviors employed in the previous experiment (i.e., voice call, web surfing, and bulk uploading). In this experiment, we place the sniffer 45m away from an eNB. We then measure the arrival signal strength by moving the UE from 10m to 150m away from the sniffer at intervals of 10m.
Figure 13 shows the measured signal strength at the attacker’s sniffer over three types of the UE’s behavior. The red box in the figure represents the detection threshold, and it was determined by measuring signal strength when uplink signals do not exist. The attacker’s sniffer easily detects the uplink signals when the UE is doing bulk uploads or making voice calls. On the other hand, the measured signal strength during web surfing is lower than the others, which causes the attacker to fail in detecting the signal when the UE is more than 60m away from the sniffer.
Appendix C Additional Considerations
C.1 Sniffer Deployment
Figure 14: Sniffer deployment in three-sectorized LTE cell.
Here we present a simple strategy in conducting UMA in the real-world environment considering practical issues. One of several sniffers used for multiangulation should be placed nearby an eNB, as a reference point. This is because if the victim is located nearby the eNB, the victim’s UE transmits uplink data with low signal power.
After AoA measurement at the reference point, the attacker can place one or two sniffers at other points based on the estimated direction at the reference point. In general, an LTE cell consists of three sector sites [6], and each sector is in charge of ${}^{\circ}$ range of radio networks (i.e., $3\times 120^{\circ}$). Thus, the attacker can place the sniffers in one sector range as shown in Figure 14. If the attacker tries three-dimensional localization, it is recommended to use three sniffers for multiangulation. Otherwise, using only two sniffers is enough since we have empirically confirmed that it shows better performance for two-dimensional localization.
Finally, when the shadow area occurs during the localization attack, the attacker can enforce the victim’s UE to increase transmission power by exploiting power boosting attack which we discuss in § 4.3.